a runner can run 4 miles in 30 minutes what is there speed in mph?
Found 2 solutions by oberobic, Alan3354:
Answer by oberobic(2304)

Their speed = 4 miles/30 min.
30 min = 1/2 hr
4/(1/2) = 8 miles/hr

Answer by Alan3354(69443)

a runner can run 4 miles in 30 minutes what is there speed in mph?
----------
There speed? Do you mean "their speed?"
Their implies plural, you said "a runner" ????
----------
Is English not your language?
------
30 minutes = 0.5 hours
d = r*t
r = d/t